### v1

# Surf-NeRF: Surface Regularised Neural Radiance Fields

Abstract
Neural Radiance Fields (NeRFs) provide a high fidelity, continuous scene representation that can realistically represent complex behaviour of light. Despite recent works like Ref-NeRF improving geometry through physics-inspired models, the ability for a NeRF to overcome shape-radiance ambiguity and converge to a representation consistent with real geometry remains limited. We demonstrate how curriculum learning of a surface light field model helps a NeRF converge towards a more geometrically accurate scene representation. We introduce four additional regularisation terms to impose geometric smoothness, consistency of normals and a separation of Lambertian and specular appearance at geometry in the scene, conforming to physical models. Our approach yields improvements of 14.4% to normals on positionally encoded NeRFs and 9.2% on grid-based models compared to current reflection-based NeRF variants. This includes a separated view-dependent appearance, conditioning a NeRF to have a geometric representation consistent with the captured scene. We demonstrate compatibility of our method with existing NeRF variants, as a key step in enabling radiance-based representations for geometry critical applications. Project page: https://roboticimaging.org/Projects/SurfNeRF
1 Introduction
Neural Radiance Fields (NeRFs) [27] provide an efficient coordinate based scene representation with wide applications to computer vision, robotics and beyond. The ray-based volumetric rendering formulation produces realistic novel views of a scene, encompassing view-dependent scene appearance. Complex appearances like specularity, transparency and interreflection lead to shape-radiance ambiguity where scene geometry is not uniquely represented, often producing imagined geometries.
Figure 1: Surf-NeRF uses the properties of surface light fields to regularise a NeRF by querying the local region of samples. These samples are used to improve consistency and smoothness in these regions accumulating density at a surface, leading to improved geometry and more physically viable appearance components (ours, right) compared to current state-of-the-art ([36], left) whilst maintaining visual fidelity. By changing how frequently regularisation occurs during training, we can control the strength of the effect.
Whilst re-parameterisation of the radiance in the scene as in Ref-NeRF [36] has shown improvements over previous state-of-the-art variants, there exists an incomplete separation of the scene’s Lambertian appearance from the specular and no explicit constraint on the placement of density. It is still common for the NeRF to place regions of density behind surfaces or in front of the camera as detached geometry (floaters) to explain complex phenomena. This leads to a non-realistic geometric scene representation, as in Figure 1.
Applications like robotics and 3D modelling require both accurate geometry and realistic appearance. The poor physical geometry currently poses a large hurdle to the widespread adoption of NeRFs as a scene representation for geometry critical tasks. Robotic manipulation and autonomous navigation and mapping require an accurate scene structure to minimise the gap between a representation and the real world, for example when grasping metal objects or navigating near reflective windows. Similarly, traditional structure-from-motion pipelines have great difficulty in reconstructing visually complex objects including reflective and transparent surfaces.
We address geometric inaccuracy arising from view-dependent phenomena with the insight that there are multiple formulations of the plenoptic function which can produce viable scene representations. A surface light field [41] describes the plenoptic function as light originating from a geometrically smooth surface. Our insight lies in that we can push the NeRF towards a more geometrically accurate representation in the same rendering framework using additional geometric and appearance based regularisation via curriculum learning. We use a first surface assumption to describe the location of geometry in the scene. Using a second sampling of the NeRF at these points, we regularise density to produce smoothly-varying normals and thin, continuous sheets of density which more realistically represent scene geometry. By enforcing view-dependent properties of a reflection light model at these points, we enable a separation of the Lambertian appearance of a scene from the view-dependent component, reducing shape-radiance ambiguity and encouraging more correct structure of the surface.
In this work, we make the following contributions:
  1. 1.
    We devise a novel regularisation approach which uses the structure of a neural radiance field to sample density, normals and appearance in the vicinity of geometry in the scene, allowing for additional representation-driven regularisation terms to be applied.
  2. 2.
    We apply local regularisation consistent with a surface light field radiance model, including geometric smoothness of density, local consistency of normals and a physically correct separation of Lambertian and specular appearance using a light interaction model.
  3. 3.
    We leverage curriculum learning of a NeRF towards a more accurate geometric scene representation which maintains visual fidelity whilst refining the density representation of the scene.
Whilst we benchmark our approach on state-of-the-art physics based NeRF variants, our methodology may also be applied to other NeRF frameworks.
This work is a key step in the deployment of NeRFs as a scene representation where both geometric and visual fidelity are critical, like robotic manipulation and navigation in complex unstructured environments.
2 Related Work
Neural Scene Representations: Neural field approaches [35, 14] to scene representation produce continuous and often high visual fidelity depictions which are able to be queried anywhere within the training set, balancing visual and geometric fidelity. In this work, we seek to improve the geometry of a NeRF whilst maintaining visual fidelity.
NeRFs [27] leverage the efficiency of ray-based construction similar to light fields with a volumetric scene representation of points which emit light. Subsequent works have improved the fidelity of representation [3, 4, 6], recovery of a static scene [32, 34, 26] and its performance around view-dependent phenomena such as reflections [36, 22]. However, relying on a purely volumetric rendering approach still allows the network to imagine geometries to explain view-dependent phenomena, particularly in cases where regions of the scene are underconstrained [31, 23, 29] or where light does not follow the physical model of a straight ray through the scene [8]. Additional regularisation terms have been shown to dramatically improve the quality of rendering in these cases [29, 23]. Improving the accuracy of normals in the scene also significantly helps learning around complex appearance [25, 36]. More recent works [37] have shown that ray-tracing multiple reflection rays can improve the representation of reflections, albeit having extremely high computation cost. In this work we use properties of surface light fields to reduce the reliance of NeRFs on additional geometries. This approach explains complex visual phenomena entirely through view-dependent appearance, whilst maintaining the formulation of a single-pass volumetric rendering approach and its quality. We achieve this using a surface light field model and local characteristics of geometry and appearance of the scene.
Surface Representations: Signed distance fields (SDFs) have garnered significant attention [44, 45, 16, 18, 2, 30] as they can represent smooth and continuous geometries in space and may be generated from multi-view constraints alone. Applying view-dependent colour channels to the SDF [38, 44, 18, 39, 40] in a similar fashion to NeRFs has demonstrated improved accuracy and continuity of representation. For diffuse objects, this representation provides a smooth and high quality reconstruction, however, more highly view-dependent appearance and complex geometry (thin structures, concave geometries) are not well represented. Some works have included a reflection parameterisation [21], however struggle to reconstruct 3D scenes with fine detail, thin structures and sharp changes compared to volumetric scenes. Appearance and geometry have an intrinsic relationship within neural representations, with degradation in geometric accuracy often resulting in altered geometry to meet appearance [18]. Our work maintains a volumetric scene representation allowing for complex geometry, but with separated view-dependent and -independent appearance. Given NeRFs are a high fidelity visual representation, we are concerned only with introducing a surface-like structure to the density field through volumetric rendering. This improves the geometry and consistency of novel view rendering by considering the NeRF as learning density constrained to a surface with smooth view-dependent terms.
Appearance Vs. Geometry: Inverse rendering seeks to produce a definite [20] separation of the scenes appearance, geometry and environment, which are necessary to create realistic renderings of scene’s under new conditions. Performing this without prior knowledge of the scene proves to be an immensely difficult task [46, 11, 15, 43], however, utilising physically-based rendering fused with learnt appearances [10, 12] provides a sufficiently constrained framework to acquire the components of the scene. Radiance field approaches seeking to learn an accurate scene representation under few view scenarios have used depth consistency [33], additional depth supervision [49] and regularisation using priors [29]. Incorporating an understanding of the physical interactions of light within a scene and its effect on appearance [47] under a solid surface assumption enables accurate geometry to be learnt alongside appearance. In the presence of specularities and other visual phenomena it is difficult to disentangle where appearance has been baked into geometry [42]. Without the need for recovering a bidirectional reflectance distribution function (BRDF), light field approaches [19] provide a more generalised framework. This enables a clear separation between appearance and geometry, by reducing the need to acquire environmental view-dependent effects. Our work leverages similar light-field characteristics to recover geometry more accurately within the volumetric rendering framework of NeRF, adding a prior to how geometry and appearance should present in the scene.
3 Method
Figure 2: An overview of the Surf-NeRF methodology. We use a first surface assumption of light to locate samples $\mathbf{x}_{\star}$ which are likely to lie on a surface in the scene. Sampling at this point in multiple directions, and at points nearby using points drawn from a sphere $\mathbf{p}$, we impose regularisation on directional and spatial behaviour in this region. We separate the Lambertian component from the specular colour channel $\mathbf{c}_{s}$ by sampling points through multiple directions. Using neighbouring points we also regularise geometric smoothness on density $\tau$ and consistency of normals $\hat{\mathbf{n}}$ leading to improved, continuous geometry.
Surf-NeRF introduces novel regularisation to locally enforce the properties of surface light fields. By regularising towards this representation, we represent smooth geometry with a physically viable separation of appearance and geometry producing more geometrically accurate radiance fields. This reduces shape radiance ambiguity by encouraging continuous regions of density with a smoothly varying view-dependent appearance. A visual depiction of our methodology is shown in Figure 2.
3.1 Preliminaries
A surface light field is a subset of the plenoptic function [9] that provides the colour of a light ray originating from a surface within the scene. Solid scene geometries are well represented given the decoupled parameterisation of the scene geometry and radiance on these surfaces.
A surface light field $L$ exists strictly on a surface geometry $G$ mapping directions in the unit 2-sphere ($\mathbf{S}^{2}$) to radiance, $L\,:\,G\times\mathbf{S}^{2}\rightarrow\mathbf{c}$, for an RGB colour triplet $\mathbf{c}$ [41].
Few real surfaces have an entirely view-dependent appearance. Similar to traditional light fields [19], a surface light field may be decomposed into a Lambertian (diffuse) reflectance, $S$, and specular (or more generally a view-dependent) component, $R$ [41, 24], $L(\mathbf{u},\boldsymbol{\omega})=S(\mathbf{u})+R(\mathbf{u},\boldsymbol{\omega})$. These intrinsic decompositions are defined for a point on the surface, $\mathbf{u}$, and viewing direction, $\boldsymbol{\omega}$. The diffuse reflectance varies only with position over the surface, whilst the view-dependent component captures elements like reflection and refraction at the surface [20].
A surface light field has a strong assumption that the radiance seen from a given direction is provided entirely from a continuous geometry in the scene. This accordingly models phenomena such as volumetric scattering, reflection or transmission as a function of surface position rather than in a volumetric (NeRF) or physics-based (rendering engine) manner. In this way, a surface light field is decoupled from the geometry of a scene, meaning a surface may be deformed while its appearance looking along a ray is maintained [41]. Our approach is motivated by this decoupling to encourage Lambertian radiance to exist on smooth sheets of density, or surfaces, with a view-dependent colour.
3.2 Model
Our proposed method builds on the reflection parameterisation in Ref-NeRF [36], which splits the scene into a diffuse and specular appearance term similar to a surface light field. This parameterisation provides enhanced results across Lambertian and specular scenes over original NeRF variants [27, 3] by learning a spatially varying diffuse colour $\mathbf{c}_{d}$, specular tint $\mathbf{s}$, rendering normal $\hat{\mathbf{n}}^{\prime}$ and a view-dependent specular colour $\mathbf{c}_{s}$ for each point in the scene. The final colour of a ray in this parameterisation is given as $\mathbf{c}=\mathbf{c}_{d}+\mathbf{s}\mathbf{c}_{s}$, a linear combination of the diffuse and specular terms. Ref-NeRF struggles to entirely separate Lambertian and specular appearance, utilising density instead to explain complex phenomena such as non-planar reflection, anisotropic reflection and interreflection, leading to the results seen in Figure 1. We apply our regularisation losses denoted by $\mathcal{L}$ to ZipNeRF [6], leveraging its state of the art performance with grid-based encoding [28], using the Ref-NeRF parameterisation. We make no modifications to the model itself beyond this, but include a second, data-driven sampling to impose physically-inspired regularisation. We maintain two proposal networks and one NeRF network with 64 and 32 samples per ray. Importantly, our regularisation terms are extensible to other NeRF variants, and may be applied after the main training as a fine-tuning stage, as we show in Section 4.
3.3 Sampling Radiance at a Surface
We regularise the surface and its light field by sampling a batch of positions and directions at the point where the ray insects a surface in the scene, as shown in Figure 3. We formulate local regularisation terms based on surface light field properties and the current scene geometry and appearance at this location.
Where prior works have sampled unseen image patches [29] to encourage consistent depths, we sample a batch of unseen rays localised at a surface point in the scene to encourage local geometric continuity. This batch approach also has significant benefit over single perturbed points seen in prior work [30, 48], as it allows for changes in surface orientation and structure not captured by a single sample to be accounted for, as we detail in Section 3.4.
We utilise a first-surface assumption to infer the location of geometry; we assume the majority of radiance is emitted by dense points closest to the camera. Our candidate surface is the first point along a ray with weight $w(\mathbf{x}_{\star})$ greater than the median weight of the ray $\mathrm{med}({\mathbf{w}}(\mathbf{x}))$, preventing regularisation from occurring behind the true location of the surface. This minimises sampling around points which are occluded and therefore not well positioned with multi-view constraints. Choosing the median ensures that this selection is more robust to skewed weight distributions, particularly early during training. This is shown in Figure 2 left.
Using the point ($\mathbf{x}_{\star}$), origin ($\mathbf{o}$), direction ($\mathbf{d}$), surface normal ($\mathbf{\hat{n}}_{\star}$), and covariance ($\mathbf{\Sigma}_{\star}$) of this sample, we generate two new batches; a spatial batch to regularise density and a directional batch to regularise appearance.
We adapt a deterministic sampling scheme on the sphere [17] to produce uniformly distributed points used in both regularisation batches. Samples drawn from a von Mises-Fischer distribution with concentration parameter $\kappa=0$ are uniformly distributed on the unit 2-sphere $\mathbf{S}^{2}\subset\mathbb{R}^{3}$. We sample this distribution deterministically using a Fibonacci-Kronecker lattice as proposed by [17] for $N$ samples. This is partitioned into $\log_{2}N$ shells sampling the unit ball. Similar to the unscented sampling presented in ZipNeRF [6], we apply a random rotation [1] to these samples during training to avoid any bias from the orientation of the sampling scheme, arriving at samples $\mathbf{p}$. Further details are provided in the supplementary material.
To produce virtual rays which can be cast during training, we use our sphere samples to define directions $\mathbf{d}_{s}=\hat{\mathbf{p}}$, origins $\mathbf{o}_{s}$ and covariances $\mathbf{\Sigma}_{s}$ to construct new conical frusta. The origins and covariances for these rays,
$\mathbf{o}_{s}=\mathbf{R}_{s}(\mathbf{o}-\mathbf{x}_{\star})+\mathbf{x}_{\star},\quad\mathbf{\Sigma}_{s}=\mathbf{R}_{s}\mathbf{\Sigma}_{\star}\mathbf{R}_{s}^{\top},$ (1a,b)
are defined by the rotation matrix $\mathbf{R}_{s}$ which rotates the original ray direction $\mathbf{d}$ to each $\mathbf{d}_{s}$. This results in $N$ rays at the same distance to the surface as the original ray, with MipNeRF gaussians [5] aligned along ray directions.
We use these new virtual rays to produce two batches, querying the density field $\tau$ in the local 3D region of the surface at $\mathbf{x}_{\star}$ and the distribution of view-dependent colours $\mathbf{c}_{s}$ through a range of viewing angles. We refer to these as the spatial and directional batches. Below we outline the locations $\mathbf{x}$, and directions $\mathbf{d}$ through which we sample the NeRF in these batches. We also visually depict our sampling schemes in Figure 3.
Figure 3: A visual depiction of the two sampling batches localised at a surface in the scene. After defining a ray surface intersection $\mathbf{x}_{\star}$, we use uniform samples on a unit ball to sample the surface through multiple viewing angles, and the local 3D volume.
Directional Sampling: Specularities at a surface are piecewise smooth, taking on the characteristics of what is being reflected. By sampling a point on a surface in a range of viewing directions and characterising the colour distribution, we can quantify how closely it matches the surface light field model introduced above. The directional batch $(\mathbf{x}_{\phi},\mathbf{d}_{\phi})$ samples at a single spatial location, but through a range of outward viewing angles:
$\quantity{\mathbf{x}_{\phi},\,\mathbf{d}_{\phi}}=\quantity{\mathbf{x}_{\star},\,\mathrm{sgn}(\mathbf{d}_{s}\cdot\mathbf{n}_{\star})\mathbf{d}_{s}},$ (2)
where $\mathrm{sgn}(\cdot)$ is the sign function.
Spatial Sampling: In a NeRF, geometry is characterised by the density field $\tau(\mathbf{x})$, whose gradient with respect to position $\mathbf{x}$ approximates surface normals $\mathbf{n}\approx-\nabla_{\mathbf{x}}\tau$. By sampling density and normals of neighbouring points, and enforcing consistency between these values we can provide additional geometric supervision towards a surface model. The spatial batch $(\mathbf{x}_{m},\mathbf{d}_{m})$ consists of points $\mathbf{p}$ in the unit ball located at the surface $\mathbf{x}_{\star}$,
$(\mathbf{x}_{m},\,\mathbf{d}_{m})=\quantity(\mathbf{x}_{\star}+\sigma_{\star,r}\mathbf{p}\sqrt{2\ln{2}},\,\mathbf{d}_{s}),$ (3)
where $\sigma_{\star,r}$ is the radial variance of the sample as per mip-NeRF [3]. This samples the NeRF within a single pixel’s conical frustum, ensuring regularisation occurs at the scale of the training data; images closer to the scene regularise at finer detail compared to those further away. Importantly, our sample volume adapts in scale towards a minimum as the NeRF localises density during training during the later stages of training. We characterise this behaviour in the supplement. As we do not care about the colour of these samples off the surface, the spatial batch uses the directions $\mathbf{d}_{s}$ of the virtual rays.
3.4 Local Smoothness
We regularise the geometry of the representation by penalising sparse and irregular density, allowing for a smooth and continuous sheet-like density field to form. These constraints are valid except where topological edges or corners exist; we therefore formulate $L_{1}$-norm regularisation terms to allow for these local features to form when required.
The spatial sampling batch encompasses the region in front of and behind the candidate surface. Using the normal at the candidate surface, we penalise points proportional to their density $\tau_{j}$ and perpendicular distance $\quantity|\frac{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}{\norm{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}}\cdot\hat{\mathbf{n}}_{\star}|$ from the surface encouraging a plane of density to form perpendicular to the normal. Points far from the surface should have low density, whilst those on the same plane as the sample point should be dense. This loss term,
$\mathcal{L}_{d}=\lambda_{d}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\quantity(1-e^{-\tau_{j}})\quantity|\frac{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}{\norm{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}}\cdot\hat{\mathbf{n}}_{\star}|,$ (4)
sums over $N$ samples in the spatial batch and is also weighted by the surface point $w_{\star}$ rendering weight. This ensures that early during training, the NeRF is able to change surface locations as more rays sample the scene. We do not place a stop gradient on $\hat{\mathbf{n}}_{\star}$, allowing for the surface to reorient based on the density of samples in the local volume.
Normals in this local region should exhibit similar smoothness, and so the term,
$\mathcal{L}_{n}=\lambda_{n}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\quantity(1-e^{-\tau_{j}})\frac{\quantity(1-\quantity(\hat{\mathbf{n}}_{j}\cdot\hat{\mathbf{n}}_{\star}))}{2},$ (5)
is imposed proportional to the angle between the normal at $\mathbf{x}_{\star}$ and those at each sample point $\mathbf{p}_{j}$. This term encourages dense samples to have a normal which is parallel with that at the surface point. Both terms localise density and encourage smooth, continuous geometries. We weight these terms by hyperparameters $\lambda_{d}$ and $\lambda_{n}$ respectively.
3.5 Visually Realistic Appearance Separation
Whilst highly specular surfaces or highly Lambertian surfaces are well represented with a reflection encoding, surfaces with both a strong Lambertian and specular component exhibit shape-radiance ambiguity. This is characterised by incorrect geometry with a view-dependent colour bias; that is the minimum specular colour over all viewing angles is non-zero to satisfy the appearance in all training view-points. More details are found in the supplementary.
To encourage the surface light field separation of the Lambertian and view-dependent terms we propose two additional losses based on directionally sampling a point. To encourage a minimum specular bias, we penalise the normalised specular colour over viewing angles in the directional sampling batch $\mathbf{c}_{s,\phi}$,
$\mathcal{L}_{b}=\lambda_{b}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\norm{\frac{\mathbf{c}_{s,j}}{\mathrm{sg}\quantity(\norm{\max_{\phi}\,\mathbf{c}_{s}})}}^{2},$ (6)
where $\mathrm{sg}(\cdot)$ is the stop gradient operator. Where the specular distribution is uniform over viewing angles, this loss term is maximal. This ensures that the appearance of a point must accumulate most of its signal in either the specular or the Lambertian term - not in both. To maintain a piecewise smooth distribution in the specular, we enforce a total variation loss over the specular colour:
$\mathcal{L}_{s}=\lambda_{s}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\mathrm{STV}_{j}(\mathbf{d}_{\phi},\mathbf{c}_{s}).$ (7)
This total variation is applied over the surface of the sampling sphere as a form of graph total variation. This term,
$\mathrm{STV}_{j}(\mathbf{d}_{\phi},\mathbf{c}_{s})=\sum_{k}\frac{1}{2}(\mathbf{d}_{j}\cdot\mathbf{d}_{k}+1)\norm{\mathbf{c}_{j}-\mathbf{c}_{k}}_{1},$ (8)
is based upon the $k$-Nearest Neighbours of each sample point and weighted by the cosine distance between the sample and its neighbours. This approach provides a localised consistency which is edge-aware continuous over the sampling sphere, capturing occlusions. Both of these terms smooth the specular distribution and increase the reliance on the diffuse component of the model and are weighted by hyperparameters $\lambda_{b}$ and $\lambda_{s}$ respectively.
3.6 Curriculum Learning
We impose our surface regularisation under curriculum learning to maintain visual fidelity whilst still improving geometry. As training progresses and additional regularisation is imposed more frequently, the representation is refined to satisfy the geometric and appearance characteristics of a surface light field. The surface constraints presented in Sections 3.4-3.5 are of similar computational complexity to a regular training step, requiring a second pass through both MLPs and therefore taking close to double the time. By using curriculum learning we trade-off the improvement to the representation with training time.
We schedule regularisation to occur in a staircase schedule of powers of 2: every 512 iterations early in the training, down to every 4 iterations in the final stages. On this schedule, there is approximately a 25% increase to training time.
For each regularisation sample, we add on each regularisation loss $\mathcal{L}$ to the losses for the combined ZipNeRF and Ref-NeRF model structure. More model implementation details are provided in the supplementary material, along with hyperparameter weights.
4 Results
Our regularisation is implemented in JAX [13] and based on the ZipNeRF codebase [7]. We utilise the improvements to quality and speed in ZipNeRF combined with the physics-based model structure introduced in Ref-NeRF. We compare to prior works MipNeRF360 [4] and Ref-NeRF [36]. We evaluate Surf-NeRF on the Shiny Real and the Shiny Objects [36] datasets used in Ref-NeRF as the main benchmark for our approach. We also evaluate on a captured dataset consisting of 4 complex reflective objects under controlled illumination, called the Koala dataset. Each scene contains approximately 40 training images and 10 test images with ground truth poses obtained using a Universal Robots UR5e robotic arm. We select objects that include non-planar specularities, fine details and specularities which present as virtual images in front of the surface. Further details are provided in the supplementary material.
4.1 Shiny Objects Dataset
Table 1 summarises results on the Shiny Objects dataset, comparing both visual fidelity by peak-signal-to-noise ratio (PSNR) in decibels (dB) and structural similarity scores (SSIM), and geometric fidelity by mean angular error (MAE) of the density derived normals in degrees and root mean squared error (RMSE) of the disparity in inverse units. We compare across positionally encoded baselines based on MipNeRF360 [4], denoted as “PE-based” and “grid-based” implementations based on ZipNeRF [6]. We also compare against a variant of MipNeRF which includes a diffuse colour channel, denoted as “MipNeRF+Diff”. We train the positionally encoded implementations using V2-8 TPUs, and the grid-based methods on four NVIDIA V100 32GB GPUs. Per scene results are included in the supplementary materials.
Table 1: Summary on the Shiny Objects [36] dataset between visual metrics (PSNR, SSIM) and geometric metrics (MAE, RMSE). Our method yields more accurate normals, trading off a maginal decrease to PSNR. Yellow, orange and red are the third, second, and first scores.
Model PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$
PE MipNeRF 31.29 0.943 56.17 0.125
[…]
Across both the positional encoding based and hash based networks we see comparable performance to Ref-NeRF in visual fidelity, demonstrating our ability to maintain visual fidelity whilst encouraging a more geometrically accurate representation. In the positionally encoded approaches we see a 14.4% improvement to normal accuracy, with notable improvements around regions of high specularity. We also see a 9.2% improvement to normals for the grid based implementation. Given the discretisation of the grid-based approaches, we see comparatively higher disparity errors across the dataset but improved normals from the interpolation of encodings. Our approach realises a 6.7% improvement in disparity versus Zip+Ref-NeRF.
In Figure 4 we demonstrate that the proposed regularisation is able to correct for warping of geometry and improve its consistency on the car scene, whilst also correctly separating the Lambertian racing stripes. Similarly, on the toaster dataset we successfully separate the Lambertian toast whilst retaining its reflection and improving geometry around specular materials with high curvature.
Figure 4: Results from the grid based models. By regularising the scene, we not only remove Lambertian scene content from the specular, like the stripes on the car and toast in the toaster, but our approach yields improved geometry particularly around curves and specularity, like the car windshield and corners of the toaster. Ref-NeRF* indicates the Zip+Ref-NeRF model.
Figure 5 depicts performance of positionally encoded models on specular surfaces with low texture. Surf-NeRF yields drastically better surface geometry and normals.
Figure 5: Geometry results on the non-grid based models. The surface regularisation can resolve errors in geometry which stem from shape-radiance ambiguity e.g. in regions with low-frequency textures on the teapot lid, without compromising visual fidelity. This attains substantially improved normals in these regions.
4.2 Shiny Real Dataset
The visual fidelity results on the Shiny Real dataset [36] are shown in Table 2 for Surf-NeRF and comparison baseline models. We demonstrate comparable visual fidelity to other state of the art NeRF variants demonstrating no appreciable drop in visual fidelity compared to non-regularised approaches on the positional encoding based models. The complex appearances of these datasets combined with the hash-encoding sees a small drop in PSNR values for grid-based variants. We show decreased reliance on floaters to explain appearance via Surf-NeRF in the supplement.
Table 2: Summary of numerical results for captured Shiny Real dataset [36]. Surf-NeRF maintains visual fidelity on captured scenes, improving SSIM over the PE based methods.
Model PSNR SSIM
[…]
MipNeRF+Diff 23.62 0.543
[…]
Surf-NeRF (Ours) 23.26 0.573
[…]
4.3 Koala Dataset
Table 3 shows the strength of our proposed regularisation on complex reflective scenes. Figure 6 demonstrates the ability for the proposed regularisation to separate Lambertian and specular colours in a more physically consistent manner, and to overcome failure cases where specularities are not densely sampled by the training data. Additional results are provided in the supplementary.
Table 3: Summary of numerical results for captured Koala dataset. Surface regularisation helps scene convergence around complex specular geometry improving PSNR.
Model PSNR SSIM
ZipNeRF 27.79 0.663
Zip+RefNeRF 23.69 0.632
Surf-NeRF (Ours) 27.24 0.655
Figure 6: The “gold head” and “shiny ball” scenes from our captured Koala dataset. The first row contains RGB and median depth renderings, whilst the second contains the diffuse and specular renderings. The Gold Head dataset shows complete separation of Lambertian and specular appearance for the mannequin. The Shiny Ball dataset represents a failure case for the Zip+Ref-NeRF baseline, however the proposed regularisation regularises these floaters away during training and appropriately places the majority of the colour into the diffuse colour term.
4.4 Ablation Study
In Table 4 we present an ablation study of our additional surface regularisation terms on the Helmet scene from the Shiny Objects dataset. As shown, including our full approach provides strong performance across both the visual and geometric metrics. Our appearance based regularisation is coupled closely to the geometry of the scene. Removing the specular bias term drastically increases the normal error, however does not alter disparity in the scene. The normal term improves continuity of the geometry in the scene. Whilst removing it allows for each individual normal to be determined more freely, reducing the MAE, it also allows for floaters to be introduced unnecessarily by our spatial term corresponding to an increase in PSNR but also an increase in disparity error within the scene.
Table 4: Ablation study on Surf-NeRF regularisation terms on the Helmet scene. All terms contribute to overall performance.
[…]
w/o spatial 32.33 0.972 10.67 0.196
[…]
Table 5: Parameter study on regularisation frequency on Coffee scene. There is a tradeoff to regularisation frequency between quality, time and effect, as in Figure 1.
Initial Freq. Final Freq. PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$
1024 8 31.38 0.967 14.89 0.181
512 4 31.99 0.968 15.29 0.181
256 2 31.70 0.964 18.17 0.208
128 1 32.09 0.967 15.95 0.181
Table 5 demonstrates an study on our curriculum learning approach. We demonstrate the effectiveness of our approach as a function of the frequency of regularisation. We demonstrate improvement to the normals and disparity in all cases compared to Zip+Ref-NeRF. Note that performance is highest for regularisation at a rate which is sparse enough for adaptation to changing density and frequent enough that corrections to the representation are retained.
4.5 Surface Regularisation as Finetuning
We demonstrate surface regularisation as a fine-tuning step, after a NeRF has been trained on pre-trained Zip+Ref-NeRF and ZipNeRF model on the car scene from the Shiny Objects dataset. As ZipNeRF has no Lambertian colour term, we do not include $\mathcal{L}_{b}$, and apply our sphere total variation regularisation $\mathcal{L}_{s}$ to the view-dependent colour $\mathbf{c}$, an example of our applicability to models without reflection parameterisation. The spatial sampling terms depend only on the geometry of the NeRF and so may be applied to other density-based neural fields without modification. We continue training at a fixed learning rate of $3.25\times 10^{-4}$ and present our results in Table 6. In both cases we demonstrate an improvement in normals over both the base model and the naïve approach where the model continues training without our regularisation. Note that finetuning with the reflection parameterisation, as in our Zip+RefNeRF model, results in higher disparity errors as the NeRF changes density placement during the additional training.
Table 6: Surf-NeRF as a finetuning step and applicability to other NeRF variants. With little additional training, we refine surface normals.
Base Extra PSNR SSIM Median MAE RMSE
ZipNeRF - 27.34 0.932 24.21 0.220
2.5k ZipNeRF 27.32 0.931 24.30 0.220
2.5k SurfNeRF 27.33 0.931 23.57 0.218
ZipRefNeRF - 30.25 0.957 18.44 0.246
2.5k ZipRefNeRF 30.14 0.956 18.59 0.251
2.5k SurfNeRF 30.13 0.956 18.29 0.248
4.6 Limitations
Modelling the scene as surfaces, elements that are better represented volumetrically like hair or subsurface scattering may be unnecessarily regularised and their representation degraded using our proposed regularisation. Since Surf-NeRF makes use of a second pass through the network, our methodology takes twice as long on each regularisation step compared to a vanilla training step. We enforce geometric consistency during training at the cost of restricting the networks ability to deliver the best photometric accuracy, leading to marginally reduced PSNR scores.
5 Conclusion
We have proposed Surf-NeRF, a novel regularisation approach and four novel physically derived regularisers, and shown these to improve geometry in a NeRF by curriculum learning towards a surface light field model. This ensures consistency of density and normals on geometry in the volume, in addition to a physics-inspired regularisation on view-dependent appearance. This enables a NeRF to adjust geometric structure during training without greatly sacrificing visual fidelity, demonstrating qualitatively improved continuity of geometry, with improvements of up to 14.4% in normals and 6.7% in disparity.
This work produces better NeRF representations for geometry critical tasks with unchanged model architectures.

### v2

@@
 Neural Radiance Fields (NeRFs) provide a high fidelity, continuous scene representation that can realistically represent complex behaviour of light.
-Despite recent works like Ref-NeRF improving geometry through physics-inspired models, the ability for a NeRF to overcome shape-radiance ambiguity and converge to a representation consistent with real geometry remains limited.
-We demonstrate how curriculum learning of a surface light field model helps a NeRF converge towards a more geometrically accurate scene representation.
-We introduce four additional regularisation terms to impose geometric smoothness, consistency of normals and a separation of Lambertian and specular appearance at geometry in the scene, conforming to physical models.
-Our approach yields improvements of 14.4% to normals on positionally encoded NeRFs and 9.2% on grid-based models compared to current reflection-based NeRF variants.
-This includes a separated view-dependent appearance, conditioning a NeRF to have a geometric representation consistent with the captured scene.
+Despite works like Ref-NeRF improving geometry through physics-inspired models, the ability for a NeRF to overcome shape-radiance ambiguity and converge to a representation consistent with real geometry remains limited.
+We demonstrate how both curriculum learning of a surface light field model and using a lattice-based hash encoding helps a NeRF converge towards a more geometrically accurate scene representation.
+We introduce four regularisation terms to impose geometric smoothness, consistency of normals, and a separation of Lambertian and specular appearance at geometry in the scene, conforming to physical models.
+Our approach yields 28% more accurate normals than traditional grid-based NeRF variants with reflection parameterisation.
+Our approach more accurately separates view-dependent appearance, conditioning a NeRF to have a geometric representation consistent with the captured scene.
 We demonstrate compatibility of our method with existing NeRF variants, as a key step in enabling radiance-based representations for geometry critical applications.
-Project page: https://roboticimaging.org/Projects/SurfNeRF
 1 Introduction
@@
 Complex appearances like specularity, transparency and interreflection lead to shape-radiance ambiguity where scene geometry is not uniquely represented, often producing imagined geometries.
-Figure 1: Surf-NeRF uses the properties of surface light fields to regularise a NeRF by querying the local region of samples.
-These samples are used to improve consistency and smoothness in these regions accumulating density at a surface, leading to improved geometry and more physically viable appearance components (ours, right) compared to current state-of-the-art ([36], left) whilst maintaining visual fidelity.
-By changing how frequently regularisation occurs during training, we can control the strength of the effect.
-Whilst re-parameterisation of the radiance in the scene as in Ref-NeRF [36] has shown improvements over previous state-of-the-art variants, there exists an incomplete separation of the scene’s Lambertian appearance from the specular and no explicit constraint on the placement of density.
+Figure 1: Surf-NeRF uses a permutohedral lattice encoding and the properties of surface light fields to regularise local regions of a NeRF.
+This improves consistency and smoothness in these regions by accumulating density at a surface, leading to improved geometry and more physically viable appearance components (ours, right) compared to current state-of-the-art ([43], left).
+Changing the frequency of regularisation balances between geometric fidelity and appearance separation.
+Re-parameterisation of the radiance in the scene as in Ref-NeRF [43] has shown improvements over previous state-of-the-art variants, however there exists an incomplete separation of the scene’s Lambertian appearance from the specular and no explicit constraint on the placement of density.
 It is still common for the NeRF to place regions of density behind surfaces or in front of the camera as detached geometry (floaters) to explain complex phenomena.
@@
 Applications like robotics and 3D modelling require both accurate geometry and realistic appearance.
-The poor physical geometry currently poses a large hurdle to the widespread adoption of NeRFs as a scene representation for geometry critical tasks.
-Robotic manipulation and autonomous navigation and mapping require an accurate scene structure to minimise the gap between a representation and the real world, for example when grasping metal objects or navigating near reflective windows.
+Poor representation of physical geometry currently poses a large hurdle to the widespread adoption of radiance-based scene representation for geometry critical tasks.
+Robotic manipulation, autonomous navigation and mapping require an accurate scene structure to minimise the gap between a representation and the real world, for example when grasping metal objects or navigating near reflective windows.
 Similarly, traditional structure-from-motion pipelines have great difficulty in reconstructing visually complex objects including reflective and transparent surfaces.
 We address geometric inaccuracy arising from view-dependent phenomena with the insight that there are multiple formulations of the plenoptic function which can produce viable scene representations.
-A surface light field [41] describes the plenoptic function as light originating from a geometrically smooth surface.
+A surface light field [48] describes the plenoptic function as light originating from a geometrically smooth surface, which we encode using a permutohedral lattice hash encoding.
 Our insight lies in that we can push the NeRF towards a more geometrically accurate representation in the same rendering framework using additional geometric and appearance based regularisation via curriculum learning.
@@
 Using a second sampling of the NeRF at these points, we regularise density to produce smoothly-varying normals and thin, continuous sheets of density which more realistically represent scene geometry.
-By enforcing view-dependent properties of a reflection light model at these points, we enable a separation of the Lambertian appearance of a scene from the view-dependent component, reducing shape-radiance ambiguity and encouraging more correct structure of the surface.
-In this work, we make the following contributions:
+Regularising the view-dependent appearance to remove Lambertian components we reduce shape-radiance ambiguity and encouraging more correct structure of the surface.
+In summary, we make the following contributions:
   1. 1.
@@
   3. 3.
-    We leverage curriculum learning of a NeRF towards a more accurate geometric scene representation which maintains visual fidelity whilst refining the density representation of the scene.
-Whilst we benchmark our approach on state-of-the-art physics based NeRF variants, our methodology may also be applied to other NeRF frameworks.
+    We introduce a permutohedral lattice encoding for NeRF which better represents non-planar and complex geometry over traditional cubic hash grids.
+  4. 4.
+    We introduce the Koala dataset of four scenes with visually challenging objects and complex geometry to benchmark our approach.
+We benchmark our approach on state-of-the-art physics based NeRF variants, however our methodology may also be applied to other NeRF frameworks and likely adapted to other radiance based representations.
 This work is a key step in the deployment of NeRFs as a scene representation where both geometric and visual fidelity are critical, like robotic manipulation and navigation in complex unstructured environments.
 2 Related Work
-Neural Scene Representations: Neural field approaches [35, 14] to scene representation produce continuous and often high visual fidelity depictions which are able to be queried anywhere within the training set, balancing visual and geometric fidelity.
-In this work, we seek to improve the geometry of a NeRF whilst maintaining visual fidelity.
+Neural Scene Representations: Implicit neural field approaches [25] produce continuous and high visual fidelity scene representations.
+These can be queried anywhere within the training set, balancing visual and geometric fidelity.
+In this work, we improve the geometry of a NeRF and maintain visual fidelity, reducing reliance on introduced geometry to meet scene appearance.
 NeRFs [34] leverage the efficiency of ray-based construction similar to light fields with a volumetric scene representation of points which emit light.
-Subsequent works have improved the fidelity of representation [3, 4, 6], recovery of a static scene [32, 34, 26] and its performance around view-dependent phenomena such as reflections [36, 22].
-However, relying on a purely volumetric rendering approach still allows the network to imagine geometries to explain view-dependent phenomena, particularly in cases where regions of the scene are underconstrained [31, 23, 29] or where light does not follow the physical model of a straight ray through the scene [8].
-Additional regularisation terms have been shown to dramatically improve the quality of rendering in these cases [29, 23].
-Improving the accuracy of normals in the scene also significantly helps learning around complex appearance [25, 36].
-More recent works [37] have shown that ray-tracing multiple reflection rays can improve the representation of reflections, albeit having extremely high computation cost.
-In this work we use properties of surface light fields to reduce the reliance of NeRFs on additional geometries.
-This approach explains complex visual phenomena entirely through view-dependent appearance, whilst maintaining the formulation of a single-pass volumetric rendering approach and its quality.
-We achieve this using a surface light field model and local characteristics of geometry and appearance of the scene.
+Subsequent works improved the fidelity of representation [4, 5, 6], recovery of a static scene [40, 42, 33] and performance around view-dependent phenomena such as reflections [43, 23].
+However, relying on a purely volumetric rendering approach allows the network to imagine geometries to explain view-dependent phenomena, particularly in cases where regions of the scene are underconstrained [38, 28, 36] or where light does not follow the physical model of a straight ray through the scene [8].
+Additional regularisation terms can dramatically improve the quality of rendering in these cases [36, 28], e.g.
+improving the accuracy of normals significantly helps representation of complex appearances [32, 43].
+More recent works [44] have shown that ray-tracing multiple reflection rays can improve representation reflective scene element, albeit with extremely high computation cost.
+In this work we use properties of surface light fields to reduce the reliance of NeRFs on additional geometries, maintaining a single-pass volumetric rendering approach.
+Radiance Based Scenes: Recent advances in explicit representations like 3D Gaussian splatting [27] (3DGS) have demonstrated improved rendering quality over volumetric implicit models, reflective objects rely heavily on introduced Gaussians behind surfaces to represent view dependence.
+Although subsequent work has sought to improve the geometric fidelity of the resulting representation [24, 26, 29, 47, 46], these involve substantial changes to the underlying representation or simplified geometry models due to the sparsity of the explicit radiance field.
+Our approach derives key benefits from the continuity of implicit radiance fields to formulate regularisation terms.
+We anticipate similarly inspired terms might be formulated for explicit methods like 3DGS, Radiant Foam and others [27, 22, 31, 24].
 Surface Representations: Signed distance fields (SDFs) have garnered significant attention [51, 52, 16, 18, 3, 37] as they can represent smooth and continuous geometries in space and may be generated from multi-view constraints alone.
-Applying view-dependent colour channels to the SDF [38, 44, 18, 39, 40] in a similar fashion to NeRFs has demonstrated improved accuracy and continuity of representation.
-For diffuse objects, this representation provides a smooth and high quality reconstruction, however, more highly view-dependent appearance and complex geometry (thin structures, concave geometries) are not well represented.
-Some works have included a reflection parameterisation [21], however struggle to reconstruct 3D scenes with fine detail, thin structures and sharp changes compared to volumetric scenes.
-Appearance and geometry have an intrinsic relationship within neural representations, with degradation in geometric accuracy often resulting in altered geometry to meet appearance [18].
-Our work maintains a volumetric scene representation allowing for complex geometry, but with separated view-dependent and -independent appearance.
-Given NeRFs are a high fidelity visual representation, we are concerned only with introducing a surface-like structure to the density field through volumetric rendering.
-This improves the geometry and consistency of novel view rendering by considering the NeRF as learning density constrained to a surface with smooth view-dependent terms.
+For diffuse planar objects, this representation provides a smooth and high-quality reconstruction; however, more highly view-dependent appearance and complex geometry (thin structures, concave geometries) are not well represented [45, 39, 21].
+Alternate volumetric encodings better suited to complex geometry like PermutoSDF [39] have substantially improved this representation, which we leverage in this work.
+Some works have included a reflection parameterisation [21], however, struggle to reconstruct 3D scenes with fine detail, thin structures, and sharp changes compared to volumetric scenes.
+Appearance and geometry have an intrinsic relationship within neural representations, with geometry being altered to account for non-Lambertian appearance [18].
+We maintain a volumetric representation allowing for complex geometry, but with separated view-dependent and -independent appearance.
 Appearance Vs.
-Geometry: Inverse rendering seeks to produce a definite [20] separation of the scenes appearance, geometry and environment, which are necessary to create realistic renderings of scene’s under new conditions.
-Performing this without prior knowledge of the scene proves to be an immensely difficult task [46, 11, 15, 43], however, utilising physically-based rendering fused with learnt appearances [10, 12] provides a sufficiently constrained framework to acquire the components of the scene.
+Geometry: Inverse rendering with mesh-based models produces a definite [20] separation of the scene’s appearance, geometry and environment, which are necessary to create realistic renderings of scenes under new conditions.
+Performing this without prior knowledge of the scene proves to be an immensely difficult task [53, 12, 14, 50], however, utilising physically-based rendering fused with learnt appearance [10, 11] provides sufficient constraints to recover the components of the scene.
 Radiance field approaches seeking to learn an accurate scene representation under few view scenarios have used depth consistency [41], additional depth supervision [56] and regularisation using priors [36].
@@
 Without the need for recovering a bidirectional reflectance distribution function (BRDF), light field approaches [19] provide a more generalised framework.
-This enables a clear separation between appearance and geometry, by reducing the need to acquire environmental view-dependent effects.
-Our work leverages similar light-field characteristics to recover geometry more accurately within the volumetric rendering framework of NeRF, adding a prior to how geometry and appearance should present in the scene.
+This enables a clear separation between appearance and geometry, reducing the need to acquire environmental view-dependent effects.
+Our work leverages similar light-field characteristics to recover geometry more accurately within the volumetric framework of NeRF, adding a prior to how geometry and appearance should present in the scene.
 3 Method
 Figure 2: An overview of the Surf-NeRF methodology.
-We use a first surface assumption of light to locate samples $\mathbf{x}_{\star}$ which are likely to lie on a surface in the scene.
-Sampling at this point in multiple directions, and at points nearby using points drawn from a sphere $\mathbf{p}$, we impose regularisation on directional and spatial behaviour in this region.
-We separate the Lambertian component from the specular colour channel $\mathbf{c}_{s}$ by sampling points through multiple directions.
+We use a first surface assumption of light to locate samples $\mathbf{x}_{\star}$ which are likely to lie on a surface in the scene during regular training steps.
+Sampling at the surface in multiple directions, and at nearby points drawn from a sphere $\mathbf{p}$, we impose regularisation on directional and spatial behaviour in this region.
+We separate the Lambertian component from the specular colour channel $\mathbf{c}_{s}$ by using a total variation prior.
 Using neighbouring points we also regularise geometric smoothness on density $\tau$ and consistency of normals $\hat{\mathbf{n}}$ leading to improved, continuous geometry.
@@
 Few real surfaces have an entirely view-dependent appearance.
-Similar to traditional light fields [19], a surface light field may be decomposed into a Lambertian (diffuse) reflectance, $S$, and specular (or more generally a view-dependent) component, $R$ [41, 24], $L(\mathbf{u},\boldsymbol{\omega})=S(\mathbf{u})+R(\mathbf{u},\boldsymbol{\omega})$.
-These intrinsic decompositions are defined for a point on the surface, $\mathbf{u}$, and viewing direction, $\boldsymbol{\omega}$.
-The diffuse reflectance varies only with position over the surface, whilst the view-dependent component captures elements like reflection and refraction at the surface [20].
-A surface light field has a strong assumption that the radiance seen from a given direction is provided entirely from a continuous geometry in the scene.
+Similar to traditional light fields [19], a surface light field may be decomposed into a Lambertian (diffuse) reflectance and a view-dependent component [48, 30].
+A surface light field has a strong assumption that the radiance seen from a given direction is provided entirely from a piecewise smooth geometry in the scene.
 This accordingly models phenomena such as volumetric scattering, reflection or transmission as a function of surface position rather than in a volumetric (NeRF) or physics-based (rendering engine) manner.
-In this way, a surface light field is decoupled from the geometry of a scene, meaning a surface may be deformed while its appearance looking along a ray is maintained [41].
+In this way, a surface light field is decoupled from the geometry of a scene, meaning a surface may be deformed while its appearance is maintained [48].
 Our approach is motivated by this decoupling to encourage Lambertian radiance to exist on smooth sheets of density, or surfaces, with a view-dependent colour.
 3.2 Model
-Our proposed method builds on the reflection parameterisation in Ref-NeRF [36], which splits the scene into a diffuse and specular appearance term similar to a surface light field.
+Reflection Parameterisation: Our method builds on the reflection parameterisation in Ref-NeRF [43], which splits the scene into a diffuse and specular appearance term similar to a surface light field.
 This parameterisation provides enhanced results across Lambertian and specular scenes over original NeRF variants [34, 4] by learning a spatially varying diffuse colour $\mathbf{c}_{d}$, specular tint $\mathbf{s}$, rendering normal $\hat{\mathbf{n}}^{\prime}$ and a view-dependent specular colour $\mathbf{c}_{s}$ for each point in the scene.
@@
 Ref-NeRF struggles to entirely separate Lambertian and specular appearance, utilising density instead to explain complex phenomena such as non-planar reflection, anisotropic reflection and interreflection, leading to the results seen in Figure 1.
-We apply our regularisation losses denoted by $\mathcal{L}$ to ZipNeRF [6], leveraging its state of the art performance with grid-based encoding [28], using the Ref-NeRF parameterisation.
-We make no modifications to the model itself beyond this, but include a second, data-driven sampling to impose physically-inspired regularisation.
+Permutohedral Lattice Hash Encoding: We replace the regular cubic hash grid used in prior NeRF variants [7, 35] with a hash-based permutohedral lattice [1].
+Prior work [39], has used this structure for its memory efficiency, but our insight lies in the volumetric coverage and relative density of simplices in this structure.
+A permutohedral lattice, is the most efficient lattice solution the sphere covering problem [15] in low dimensions.
+Vertices in this lattice are arranged evenly throughout $d$-space and packed more densely [1], allowing for more accurate interpolation of features over cubic volume elements.
+In $d=3$-space, simplices take the form of regular tetrahedra which can more accurately represent complex non-planar geometries and closely approximating the true geometry of the scene.
+Our permutohedral lattice encoding therefore has two main benefits over cubic grids: 1) half the memory accesses per interpolation by using barycentric interpolation of a tetrahedron [39], and 2) dramatically improved representation of non-planar surfaces derived from density by representing curved geometry with multi-resolution tetrahedral features.
+We apply our regularisation losses denoted by $\mathcal{L}$ and permutohedral lattice encoding to ZipNeRF [6] using the Ref-NeRF parameterisation, leveraging its state of the art performance with grid-based encoding [35].
+We make no modifications to the model itself beyond our hash encoding, but include a second, data-driven sampling to impose physically-inspired regularisation.
 We maintain two proposal networks and one NeRF network with 64 and 32 samples per ray.
-Importantly, our regularisation terms are extensible to other NeRF variants, and may be applied after the main training as a fine-tuning stage, as we show in Section 4.
+Importantly, our regularisation terms are extensible to other NeRF variants, and may be applied after the main training as a fine-tuning stage.
+An overview of model structure is provided in the supplementary materials.
 3.3 Sampling Radiance at a Surface
-We regularise the surface and its light field by sampling a batch of positions and directions at the point where the ray insects a surface in the scene, as shown in Figure 3.
-We formulate local regularisation terms based on surface light field properties and the current scene geometry and appearance at this location.
+We regularise the surface and its light field by sampling a batch of positions and directions at the point where the ray intersects a surface in the scene, as shown in Figure 3.
+These local regularisation terms are based on surface light field properties, current scene geometry and appearance at this location.
 Where prior works have sampled unseen image patches [36] to encourage consistent depths, we sample a batch of unseen rays localised at a surface point in the scene to encourage local geometric continuity.
 This batch approach also has significant benefit over single perturbed points seen in prior work [37, 55], as it allows for changes in surface orientation and structure not captured by a single sample to be accounted for, as we detail in Section 3.4.
-We utilise a first-surface assumption to infer the location of geometry; we assume the majority of radiance is emitted by dense points closest to the camera.
-Our candidate surface is the first point along a ray with weight $w(\mathbf{x}_{\star})$ greater than the median weight of the ray $\mathrm{med}({\mathbf{w}}(\mathbf{x}))$, preventing regularisation from occurring behind the true location of the surface.
+We utilise a first-surface assumption to infer the location of geometry; we assume the majority of radiance is emitted by the point of highest density closest to the camera.
+Our candidate surface is the first point along a ray with weight greater than the median weight of the ray, preventing regularisation from occurring behind the true location of the surface.
 This minimises sampling around points which are occluded and therefore not well positioned with multi-view constraints.
@@
 This is shown in Figure 2 left.
-Using the point ($\mathbf{x}_{\star}$), origin ($\mathbf{o}$), direction ($\mathbf{d}$), surface normal ($\mathbf{\hat{n}}_{\star}$), and covariance ($\mathbf{\Sigma}_{\star}$) of this sample, we generate two new batches; a spatial batch to regularise density and a directional batch to regularise appearance.
+Using the point $\mathbf{x}_{\star}$, ray origin $\mathbf{o}$ and direction $\mathbf{d}$, surface normal $\mathbf{\hat{n}}_{\star}$, and covariance $\mathbf{\Sigma}_{\star}$ of this sample’s integrated positional encoding [4], we generate two new batches: a spatial batch to regularise density and a directional batch to regularise appearance.
+Spatial samples query the density field $\tau$ in the local 3D region of the surface at $\mathbf{x}_{\star}$ and the directional samples query the distribution of view-dependent colour $\mathbf{c}_{s}$ at different viewing angles.
 We adapt a deterministic sampling scheme on the sphere [17] to produce uniformly distributed points used in both regularisation batches.
@@
 Further details are provided in the supplementary material.
-To produce virtual rays which can be cast during training, we use our sphere samples to define directions $\mathbf{d}_{s}=\hat{\mathbf{p}}$, origins $\mathbf{o}_{s}$ and covariances $\mathbf{\Sigma}_{s}$ to construct new conical frusta.
+We construct both batches of samples from virtual rays passing through this sphere centred at the surface.
+The points $p$ define directions $\mathbf{d}_{s}=\hat{\mathbf{p}}$, origins $\mathbf{o}_{s}$ and covariances $\mathbf{\Sigma}_{s}$ to construct new conical frusta.
 The origins and covariances for these rays,
@@
 This results in $N$ rays at the same distance to the surface as the original ray, with MipNeRF Gaussians [5] aligned along ray directions.
-We use these new virtual rays to produce two batches, querying the density field $\tau$ in the local 3D region of the surface at $\mathbf{x}_{\star}$ and the distribution of view-dependent colours $\mathbf{c}_{s}$ through a range of viewing angles.
-We refer to these as the spatial and directional batches.
-Below we outline the locations $\mathbf{x}$, and directions $\mathbf{d}$ through which we sample the NeRF in these batches.
-We also visually depict our sampling schemes in Figure 3.
+Using these virtual rays, we construct sampling batches at positions $\mathbf{x}$, and directions $\mathbf{d}$ to employ surface light field regularisation.
+Both batches inherit covariances from the virtual rays $\mathbf{\Sigma}_{s}$ for anti-aliasing as in ZipNeRF [7].
 Figure 3: A visual depiction of the two sampling batches localised at a surface in the scene.
@@
 The directional batch $(\mathbf{x}_{\phi},\mathbf{d}_{\phi})$ samples at a single spatial location, but through a range of outward viewing angles:
-$\quantity{\mathbf{x}_{\phi},\,\mathbf{d}_{\phi}}=\quantity{\mathbf{x}_{\star},\,\mathrm{sgn}(\mathbf{d}_{s}\cdot\mathbf{n}_{\star})\mathbf{d}_{s}},$ (2)
+$\quantity(\mathbf{x}_{\phi},\,\mathbf{d}_{\phi})=\quantity(\mathbf{x}_{\star},\,\mathrm{sgn}(\mathbf{d}_{s}\cdot\mathbf{n}_{\star})\mathbf{d}_{s}),$ (2)
 where $\mathrm{sgn}(\cdot)$ is the sign function.
@@
 where $\sigma_{\star,r}$ is the radial variance of the sample as per mip-NeRF [4].
-This samples the NeRF within a single pixel’s conical frustum, ensuring regularisation occurs at the scale of the training data; images closer to the scene regularise at finer detail compared to those further away.
-Importantly, our sample volume adapts in scale towards a minimum as the NeRF localises density during training during the later stages of training.
-We characterise this behaviour in the supplement.
-As we do not care about the colour of these samples off the surface, the spatial batch uses the directions $\mathbf{d}_{s}$ of the virtual rays.
+This samples the NeRF within a single pixel’s conical frustum, ensuring regularisation occurs at the scale of the training data; images closer to the scene regularise at finer detail.
+Our sample volume adapts in scale towards a minimum as the NeRF localises density during training during the later stages of training, as shown in the supplement.
+Since the colours go unused, the spatial batch uses the directions $\mathbf{d}_{s}$ of the virtual rays.
 3.4 Local Smoothness
@@
 The spatial sampling batch encompasses the region in front of and behind the candidate surface.
-Using the normal at the candidate surface, we penalise points proportional to their density $\tau_{j}$ and perpendicular distance $\quantity|\frac{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}{\norm{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}}\cdot\hat{\mathbf{n}}_{\star}|$ from the surface encouraging a plane of density to form perpendicular to the normal.
-Points far from the surface should have low density, whilst those on the same plane as the sample point should be dense.
+Using the normal at the candidate surface, we penalise points proportional to their density $\tau_{j}$ and perpendicular distance from the surface encouraging a plane of density to form perpendicular to the normal.
+Points far from the surface should have low density, and those on the same plane as the sample point should be dense.
 This loss term,
@@
 3.5 Visually Realistic Appearance Separation
-Whilst highly specular surfaces or highly Lambertian surfaces are well represented with a reflection encoding, surfaces with both a strong Lambertian and specular component exhibit shape-radiance ambiguity.
+Although highly specular surfaces or highly Lambertian surfaces are well represented with a reflection encoding, surfaces with both a strong Lambertian and specular component exhibit shape-radiance ambiguity.
 This is characterised by incorrect geometry with a view-dependent colour bias; that is the minimum specular colour over all viewing angles is non-zero to satisfy the appearance in all training view-points.
@@
 This approach provides a localised consistency which is edge-aware continuous over the sampling sphere, capturing occlusions.
-Both of these terms smooth the specular distribution and increase the reliance on the diffuse component of the model and are weighted by hyperparameters $\lambda_{b}$ and $\lambda_{s}$ respectively.
+Both of these terms smooth the specular distribution and increase the reliance on the diffuse component of the model.
+These are weighted by hyperparameters $\lambda_{b}$ and $\lambda_{s}$ respectively.
 3.6 Curriculum Learning
-We impose our surface regularisation under curriculum learning to maintain visual fidelity whilst still improving geometry.
+We impose our surface regularisation under curriculum learning to maintain visual fidelity and improve geometry.
 As training progresses and additional regularisation is imposed more frequently, the representation is refined to satisfy the geometric and appearance characteristics of a surface light field.
 The surface constraints presented in Sections 3.4-3.5 are of similar computational complexity to a regular training step, requiring a second pass through both MLPs and therefore taking close to double the time.
-By using curriculum learning we trade-off the improvement to the representation with training time.
+By using curriculum learning we trade off the improvement to the representation with training time.
 We schedule regularisation to occur in a staircase schedule of powers of 2: every 512 iterations early in the training, down to every 4 iterations in the final stages.
@@
 We evaluate Surf-NeRF on the Shiny Real and the Shiny Objects [43] datasets used in Ref-NeRF as the main benchmark for our approach.
-We also evaluate on a captured dataset consisting of 4 complex reflective objects under controlled illumination, called the Koala dataset.
+We also introduce a captured dataset consisting of 4 complex reflective objects under controlled illumination, called the Koala dataset.
 Each scene contains approximately 40 training images and 10 test images with ground truth poses obtained using a Universal Robots UR5e robotic arm.
@@
 4.1 Shiny Objects Dataset
-Table 1 summarises results on the Shiny Objects dataset, comparing both visual fidelity by peak-signal-to-noise ratio (PSNR) in decibels (dB) and structural similarity scores (SSIM), and geometric fidelity by mean angular error (MAE) of the density derived normals in degrees and root mean squared error (RMSE) of the disparity in inverse units.
-We compare across positionally encoded baselines based on MipNeRF360 [4], denoted as “PE-based” and “grid-based” implementations based on ZipNeRF [6].
-We also compare against a variant of MipNeRF which includes a diffuse colour channel, denoted as “MipNeRF+Diff”.
-We train the positionally encoded implementations using V2-8 TPUs, and the grid-based methods on four NVIDIA V100 32GB GPUs.
-Per scene results are included in the supplementary materials.
+Table 1 summarises results on the Shiny Objects dataset.
+We compare visual fidelity by peak-signal-to-noise ratio (PSNR) in decibels (dB) and structural similarity scores (SSIM), and geometric fidelity by mean angular error (MAE) of surface normals in degrees and root mean squared error (RMSE) of the disparity in inverse units.
+We compare across positionally encoded methods (PE) based on MipNeRF360 [5] and “grid-based” methods (Grid) based on ZipNeRF [6].
+This includes baselines: MipNeRF with a diffuse channel (MipNeRF+Diff), and lattice and cubic grid based ZipNeRF with reflection parameterisation (Zip+Ref-NeRF).
+We train the PE implementations using V2-8 TPUs, and the Grid methods on four NVIDIA V100 32GB GPUs.
+Scene results are included in the supplementary materials.
 Table 1: Summary on the Shiny Objects [43] dataset between visual metrics (PSNR, SSIM) and geometric metrics (MAE, RMSE).
-Our method yields more accurate normals, trading off a maginal decrease to PSNR.
+Our method yields significantly more accurate normals, trading off a marginal decrease to PSNR.
 Yellow, orange and red are the third, second, and first scores.
@@
 PE MipNeRF 31.29 0.943 56.17 0.125
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Across both the positional encoding based and hash based networks we see comparable performance to Ref-NeRF in visual fidelity, demonstrating our ability to maintain visual fidelity whilst encouraging a more geometrically accurate representation.
+MipNeRF+Diff 30.84 0.936 60.00 0.122
+Ref-NeRF 33.21 0.971 25.89 0.117
+Surf-NeRF (Reg. Only) 33.01 0.967 22.15 0.117
+Grid ZipNeRF 31.02 0.952 19.47 0.209
+Zip+Ref-NeRF 33.14 0.964 14.70 0.195
+Lattice+Ref-NeRF 33.21 0.964 13.76 0.202
+Surf-NeRF (No Lattice) 32.13 0.954 13.35 0.202
+Surf-NeRF (Ours) 32.82 0.955 10.60 0.198
+Across both the positional encoding based and hash based networks we see comparable performance to Ref-NeRF in visual fidelity, demonstrating our ability to maintain visual fidelity and encourage a more geometrically accurate representation.
 In the positionally encoded approaches we see a 14.4% improvement to normal accuracy, with notable improvements around regions of high specularity.
-We also see a 9.2% improvement to normals for the grid based implementation.
+Surf-NeRF attains a 27.9% improvement to normals for the grid based implementation.
 Given the discretisation of the grid-based approaches, we see comparatively higher disparity errors across the dataset but improved normals from the interpolation of encodings.
-Our approach realises a 6.7% improvement in disparity versus Zip+Ref-NeRF.
-In Figure 4 we demonstrate that the proposed regularisation is able to correct for warping of geometry and improve its consistency on the car scene, whilst also correctly separating the Lambertian racing stripes.
-Similarly, on the toaster dataset we successfully separate the Lambertian toast whilst retaining its reflection and improving geometry around specular materials with high curvature.
-Figure 4: Results from the grid based models.
-By regularising the scene, we not only remove Lambertian scene content from the specular, like the stripes on the car and toast in the toaster, but our approach yields improved geometry particularly around curves and specularity, like the car windshield and corners of the toaster.
+In Figure 4 we demonstrate that the proposed regularisation is able to substantially improve surface normals and the surface consistency.
+This also separates the Lambertian scene content, including the racing stripes as in 5.
+Similarly, on the toaster scene we successfully separate the Lambertian toast and retain its reflection in the specular term.
+Figure 4: Geometry comparison between normals achieved by Surf-NeRF and traditional grid implementations on the car and coffee scenes.
 Ref-NeRF* indicates the Zip+Ref-NeRF model.
-Figure 5 depicts performance of positionally encoded models on specular surfaces with low texture. Surf-NeRF yields drastically better surface geometry and normals.
-Figure 5: Geometry results on the non-grid based models.
-The surface regularisation can resolve errors in geometry which stem from shape-radiance ambiguity e.g.
-in regions with low-frequency textures on the teapot lid, without compromising visual fidelity.
-This attains substantially improved normals in these regions.
+Regularising the density field improves the continuity of surfaces reducing holes and divots.
+The inclusion of the permutohedral lattice encoding significantly improves surface normals by better approximating smooth surface curvature.
+Figure 5: Appearance separation between Surf-NeRF and baseline method Zip+Ref-NeRF.
+Colour renderings are comparable, but Surf-NeRF produces physically based separation of Lambertian and specular scene content, such the racing stripes and toast.
 4.2 Shiny Real Dataset
@@
 We demonstrate comparable visual fidelity to other state of the art NeRF variants demonstrating no appreciable drop in visual fidelity compared to non-regularised approaches on the positional encoding based models.
-The complex appearances of these datasets combined with the hash-encoding sees a small drop in PSNR values for grid-based variants.
-We show decreased reliance on floaters to explain appearance via Surf-NeRF in the supplement.
+The complex appearances of these datasets combined with the hash encoding sees a small drop in PSNR values for grid-based variants.
+Surf-NeRF shows decreased reliance on floaters to explain appearance as shown in the supplement.
 Table 2: Summary of numerical results for captured Shiny Real dataset [43]. Surf-NeRF maintains visual fidelity on captured scenes, improving SSIM over the PE based methods.
 Model PSNR SSIM
-[…]
+PE MipNeRF 23.72 0.543
 MipNeRF+Diff 23.62 0.543
-[…]
-Surf-NeRF (Ours) 23.26 0.573
-[…]
-[…]
-[…]
+Ref-NeRF 23.65 0.529
+Surf-NeRF (Reg. Only) 23.26 0.573
+Grid ZipNeRF 23.63 0.626
+Zip+RefNeRF 22.63 0.626
+Lattice+RefNeRF 22.94 0.614
+Surf-NeRF (No Lattice) 22.48 0.600
+Surf-NeRF (Ours) 23.04 0.600
 4.3 Koala Dataset
-Table 3 shows the strength of our proposed regularisation on complex reflective scenes.
+Table 3 shows the strength of the proposed regularisation on complex reflective scenes.
 Figure 6 demonstrates the ability for the proposed regularisation to separate Lambertian and specular colours in a more physically consistent manner, and to overcome failure cases where specularities are not densely sampled by the training data.
@@
 Zip+RefNeRF 23.69 0.632
-Surf-NeRF (Ours) 27.24 0.655
-Figure 6: The “gold head” and “shiny ball” scenes from our captured Koala dataset.
-The first row contains RGB and median depth renderings, whilst the second contains the diffuse and specular renderings.
-The Gold Head dataset shows complete separation of Lambertian and specular appearance for the mannequin.
-The Shiny Ball dataset represents a failure case for the Zip+Ref-NeRF baseline, however the proposed regularisation regularises these floaters away during training and appropriately places the majority of the colour into the diffuse colour term.
+Surf-NeRF (No Lattice) 27.24 0.655
+Surf-NeRF (Ours) 26.86 0.654
+Figure 6: The “Gold Head” and “Shiny Ball” scenes from the Koala dataset.
+Top rows contain RGB and median depth renderings.
+The Gold Head scene shows complete separation of Lambertian and specular appearance for the mannequin.
+The Shiny Ball dataset represents a failure case for Zip+Ref-NeRF baseline, however the proposed regularisation successfully reconstructs and appropriately places the majority of the colour into the diffuse channel.
 4.4 Ablation Study
-In Table 4 we present an ablation study of our additional surface regularisation terms on the Helmet scene from the Shiny Objects dataset.
-As shown, including our full approach provides strong performance across both the visual and geometric metrics.
+As shown in Table 4 including our full approach provides strong performance across both the visual and geometric metrics on Shiny Objects [43].
 Our appearance based regularisation is coupled closely to the geometry of the scene.
-Removing the specular bias term drastically increases the normal error, however does not alter disparity in the scene.
-The normal term improves continuity of the geometry in the scene.
-Whilst removing it allows for each individual normal to be determined more freely, reducing the MAE, it also allows for floaters to be introduced unnecessarily by our spatial term corresponding to an increase in PSNR but also an increase in disparity error within the scene.
-Table 4: Ablation study on Surf-NeRF regularisation terms on the Helmet scene. All terms contribute to overall performance.
+The permutohedral lattice encoding significantly improves surface normals coupled with our regularisation.
+Table 4: Ablation study on Surf-NeRF regularisation terms on the Shiny Objects dataset.
+All terms contribute to overall performance.
+$\bigstar$ ablates deterministic sphere sampling, # ablates the permutohedral lattice and all $\mathcal{L}$ terms match those in the paper.
 […]
-[…]
-w/o spatial 32.33 0.972 10.67 0.196
-[…]
-[…]
-[…]
-Table 5: Parameter study on regularisation frequency on Coffee scene. There is a tradeoff to regularisation frequency between quality, time and effect, as in Figure 1.
+33.03 0.962 15.47 0.181
+32.55 0.959 15.81 0.192
+32.83 0.961 14.83 0.200
+33.32 0.965 16.34 0.196
+33.12 0.964 12.21 0.183
+32.76 0.958 11.40 0.196
+32.96 0.964 10.90 0.199
+33.14 0.969 10.74 0.201
+Rand. Samples 32.26 0.954 13.70 0.203
+(No Lattice) 32.13 0.954 13.35 0.202
+(Ours) 32.82 0.955 10.60 0.198
+Table 5: Parameter study on regularisation frequency.
+There is a tradeoff to regularisation frequency between quality, time and effect, as in Figure 1.
+The 512/4 regime balanced multi-view photometric losses with the surface light field regularisation.
 Initial Freq. Final Freq. PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$
-1024 8 31.38 0.967 14.89 0.181
-512 4 31.99 0.968 15.29 0.181
-256 2 31.70 0.964 18.17 0.208
-128 1 32.09 0.967 15.95 0.181
-Table 5 demonstrates an study on our curriculum learning approach.
-We demonstrate the effectiveness of our approach as a function of the frequency of regularisation.
-We demonstrate improvement to the normals and disparity in all cases compared to Zip+Ref-NeRF.
-Note that performance is highest for regularisation at a rate which is sparse enough for adaptation to changing density and frequent enough that corrections to the representation are retained.
+1024 8 32.63 0.955 12.46 0.204
+512 4 32.82 0.953 10.60 0.198
+256 2 30.59 0.946 18.99 0.220
+128 1 29.77 0.926 28.37 0.218
+Table 5 demonstrates a parameter study on our curriculum learning approach.
+Regularising frequency forces the NeRF to learn better separation of Lambertian and view-dependent channels, as in Figure 1 but reduces overall geometric quality by strongly manipulating density earlier in training.
+See the supplementary for more detail.
 4.5 Surface Regularisation as Finetuning
 We demonstrate surface regularisation as a fine-tuning step, after a NeRF has been trained on pre-trained Zip+Ref-NeRF and ZipNeRF model on the car scene from the Shiny Objects dataset.
-As ZipNeRF has no Lambertian colour term, we do not include $\mathcal{L}_{b}$, and apply our sphere total variation regularisation $\mathcal{L}_{s}$ to the view-dependent colour $\mathbf{c}$, an example of our applicability to models without reflection parameterisation.
-The spatial sampling terms depend only on the geometry of the NeRF and so may be applied to other density-based neural fields without modification.
+Since ZipNeRF has no Lambertian colour term, we do not include $\mathcal{L}_{b}$, and apply our sphere total variation regularisation $\mathcal{L}_{s}$ to the view-dependent colour $\mathbf{c}$, demonstrating applicability to models without reflection parameterisation.
+The spatial sampling terms depend only on the geometry and can be applied to other density-based neural fields without modification.
 We continue training at a fixed learning rate of $3.25\times 10^{-4}$ and present our results in Table 6.
-In both cases we demonstrate an improvement in normals over both the base model and the naïve approach where the model continues training without our regularisation.
-Note that finetuning with the reflection parameterisation, as in our Zip+RefNeRF model, results in higher disparity errors as the NeRF changes density placement during the additional training.
-Table 6: Surf-NeRF as a finetuning step and applicability to other NeRF variants. With little additional training, we refine surface normals.
+Normals are improved over the base model and naïve approach where the model continues training without our regularisation.
+Finetuning with the reflection parameterisation, as in our Zip+RefNeRF model slightly increases disparity errors as the NeRF changes density placement during the additional training.
+Table 6: Surf-NeRF as a finetuning step with applicability to other NeRF variants. Additional training refines surface normals.
 Base Extra PSNR SSIM Median MAE RMSE
@@
 4.6 Limitations
-Modelling the scene as surfaces, elements that are better represented volumetrically like hair or subsurface scattering may be unnecessarily regularised and their representation degraded using our proposed regularisation.
-Since Surf-NeRF makes use of a second pass through the network, our methodology takes twice as long on each regularisation step compared to a vanilla training step.
-We enforce geometric consistency during training at the cost of restricting the networks ability to deliver the best photometric accuracy, leading to marginally reduced PSNR scores.
+Modelling the scene as surfaces, volumetric elements like hair or subsurface scattering may have their representation degraded with the proposed regularisation.
+Our methodology takes twice as long on each regularisation step compared to a vanilla training step owing to two passes through the network.
+We enforce geometric consistency during training at the cost of reducing photometric accuracy, leading to slightly reduced PSNR scores.
 5 Conclusion
-We have proposed Surf-NeRF, a novel regularisation approach and four novel physically derived regularisers, and shown these to improve geometry in a NeRF by curriculum learning towards a surface light field model.
-This ensures consistency of density and normals on geometry in the volume, in addition to a physics-inspired regularisation on view-dependent appearance.
-This enables a NeRF to adjust geometric structure during training without greatly sacrificing visual fidelity, demonstrating qualitatively improved continuity of geometry, with improvements of up to 14.4% in normals and 6.7% in disparity.
-This work produces better NeRF representations for geometry critical tasks with unchanged model architectures.
+We have proposed Surf-NeRF, a novel regularisation approach with four physically derived regularisation terms and a lattice hash encoding, and shown these to improve geometry in a NeRF by curriculum learning towards a surface light field model.
+This ensures smoothly varying surface normals and geometry with more realistic Lambertian and view-dependent appearance.
+This enables a NeRF to adjust geometric structure without greatly sacrificing visual fidelity, improving accuracy of up to 27.8% in normals and 6.7% in disparity.
+This work is a key step towards better radiance-based representations for geometry critical tasks.